# The inner structure and thermodynamics of a thin accretion disc

### v1

@@ 4.4 Other Considerations
The electron-only cooling function presented in this paper is by necessity a simplification of the “true” physics of a two-temperature accretion disc. This model does not include non-local effects due to Compton scattering, which could affect the truncation structure (Esin et al., 1997). In addition, it assumes that the main coupling mechanism between protons and electrons is due to Coulomb collisions. While this assumption likely holds for the conditions in an accretion disc, it is in principle possible for other coupling mechanisms to exist (Begelman & Chiueh, 1988). The form of the cooling function assumes a cooling time proportional to the orbital time, whereas the physical cooling time is likely much shorter, resulting in an electron temperature that deviates from the target temperature within the ISCO. The cooling function also by nature assumes optically-thin radiation, since the radiation energy is simply removed from the gas without interacting with any gas along the line-of-sight. Presumably this assumption will break down at higher accretion rates where the disc becomes dominated by radiation pressure.
@@ Appendix A Two-temperature thermalization tests
We show the results from three typical simulations in Figure 12. All three simulations start with an electron temperature $T_{e,0}=10^{8}$ K and a proton temperature $T_{p,0}=10^{10}$ K. In the first simulation, $n_{0}=6\times 10^{21}{\rm particles/cm^{3}}$ and $t_{\rm cool}=10^{2}~{}L/c$. With these parameters, the electrons reach the target cooling temperature before the electrons and protons thermalize. Because the cooling time is short, the electrons stay glued to the target temperature. The second simulation has a cooling time ten times that of the first, allowing the electron temperature to temporarily overshoot the target temperature. In the last simulation, a density ten times that of the first two simulations means that Coulomb collisions are much stronger initially. Coulomb collisions therefore force the gas and electron fluid to thermalize much faster. Because of the stronger Coulomb collisions, the electron temperature briefly exceeds the target temperature despite a shorter cooling time than the second simulation.

### v2

@@ 4.5 Numerical Considerations
The electron-only cooling function presented in this paper is by necessity a simplification of the “true” physics of a two-temperature accretion disc. This model does not include non-local effects due to Compton scattering, which could affect the truncation structure (Esin, 1997). In addition, it assumes that the main coupling mechanism between protons and electrons is due to Coulomb collisions. While this assumption likely holds for the conditions in an accretion disc, it is in principle possible for other coupling mechanisms to exist (Begelman & Chiueh, 1988). The form of the cooling function assumes a cooling time proportional to the orbital time, whereas the physical cooling time is likely much shorter, resulting in an electron temperature that deviates from the target temperature within the ISCO.
The cooling function also by nature assumes optically-thin radiation, since the radiation energy is simply removed from the gas without interacting with any gas along the line-of-sight. Presumably this assumption will break down at higher accretion rates where the disc becomes dominated by radiation pressure and in optically-thick regions. Because radiation pressure changes the disc temperature and density profiles, future work will address disc truncation in the radiation pressure dominated regime with a constant $H/r$ rather than a constant target temperature, as was motivated by considering Compton scattering as the dominant cooling mechanism. More accurate treatment of radiation remains computationally expensive, and requires circumventing the thermal instability, whose impact on global disks remains unclear (Mishra et al., 2016; Fragile et al., 2018; Mishra et al., 2022). The numerically tractable choice of a cooling function provides a minimal model for capturing the effects of radiation on a stable disc and permits studying the impact of radiation as a function of accretion rate.
@@ Appendix A Two-temperature thermalization tests
We show the results from three typical simulations in Figure 13. All three simulations start with an electron temperature $T_{e,0}=10^{8}$ K and a proton temperature $T_{p,0}=10^{10}$ K. In the first simulation, $n_{0}=6\times 10^{21}{\rm particles/cm^{3}}$ and $t_{\rm cool}=10^{2}~{}L/c$. With these parameters, the electrons reach the target cooling temperature before the electrons and protons thermalize. Because the cooling time is short, the electrons stay glued to the target temperature. The second simulation has a cooling time ten times that of the first, allowing the electron temperature to temporarily overshoot the target temperature. In the last simulation, a density ten times that of the first two simulations means that Coulomb collisions are much stronger initially. Coulomb collisions therefore force the gas and electron fluid to thermalize much faster. Because of the stronger Coulomb collisions, the electron temperature briefly exceeds the target temperature despite a shorter cooling time than the second simulation. See Naethe Motta et al. (2025) for a detailed comparison of numerical techniques.